OpenStudy (anonymous):

Sue plans to paste two ribbons along the diagonals of the top surface of a rectangular gift box, as shown below.

OpenStudy (anonymous):

What is the minimum length of ribbon that Sue would require?

OpenStudy (anonymous):

@jim_thompson5910 @thomaster

OpenStudy (campbell_st):

use pythagoras' theorem to find the length of 1 diagonal. In an rectangle diagonals are equal in length. So double your answer

jimthompson5910 (jim_thompson5910):

Hint:  6^2 + 8^2 = x^2

OpenStudy (anonymous):

100=x^2 x=10 double it so it's 20?

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

Thank you.

jimthompson5910 (jim_thompson5910):

yw